# GreedyML: A Parallel Algorithm for Maximizing Submodular Functions

### v1

@@ Abstract
We describe a parallel approximation algorithm for maximizing monotone submodular functions subject to hereditary constraints on distributed memory multiprocessors. Our work is motivated by the need to solve submodular optimization problems on massive data sets, for practical applications in areas such as data summarization, machine learning, and graph sparsification.
Our work builds on the randomized distributed RandGreedi algorithm, proposed by Barbosa, Ene, Nguyen, and Ward (2015). This algorithm computes a distributed solution by randomly partitioning the data among all the processors and then employing a single accumulation step in which all processors send their partial solutions to one processor. However, for large problems, the accumulation step could exceed the memory available on a processor, and the processor which performs the accumulation could become a computational bottleneck.
Here, we propose a generalization of the RandGreedi algorithm that employs multiple accumulation steps to reduce the memory required. We analyze the approximation ratio and the time complexity of the algorithm (in the BSP model). We evaluate the new GreedyML algorithm on three classes of problems, and report results from massive data sets with millions of elements. The results show that the GreedyML algorithm can solve problems where the sequential Greedy and distributed RandGreedi algorithms fail due to memory constraints. For certain computationally intensive problems, the GreedyML algorithm can be faster than the RandGreedi algorithm. The observed approximation quality of the solutions computed by the GreedyML algorithm closely matches those obtained by the RandGreedi  algorithm on these problems.
@@ 1 Introduction
We describe a scalable parallel approximation algorithm for maximizing monotone submodular functions subject to hereditary constraints on distributed memory multiprocessors. We build on an earlier distributed approximation algorithm which has limited parallelism and higher memory requirements. Although this problem is NP-hard (the objective function is nonlinear), a Greedy algorithm that maximizes the marginal gain (defined later) at each step is $(1-1/\mathrm{e})\approx 0.63$-approximate for cardinality constraints and $1/2$-approximate for matroid constraints; here $\mathrm{e}$ is Euler’s number.
@@
Combinatorial optimization with a submodular objective function (rather than a linear objective function) leads to diversity in the computed solution, since at each step the algorithm chooses an element with the least properties in common with the current solution set. A broad collection of optimization problems could be modeled using submodular functions, including data and document summarization [24], load balancing parallel computations in quantum chemistry [9], sensor selection [6], resource allocation [27], active learning [11], interpretability of neural networks [7], influence maximization in social networks [13], diverse recommendation [5] etc. Submodular optimization problems often have efficient approximation algorithms to solve them, since submodular functions have properties that make them discrete analogs of both convex and concave continuous functions. Surveys discussing submodular optimization formulations, algorithms, and computational experiments include Tohidi et al. [28] and Krause and Golovin [14].
@@
Our algorithm builds on the RandGreedi  framework [2], a state-of-the-art randomized distributed algorithm for monotone submodular function maximization under hereditary constraints, which has an approximation ratio half that of the Greedy algorithm. The RandGreedi algorithm randomly partitions the data among all the processors, runs the standard Greedy algorithm on each partition independently in parallel, and then executes a single accumulation step in which all processors send their partial solutions to one processor. However, this step could exceed the memory available on a processor when the memory is small relative to the size of the data, or when solutions are large. Additionally, this merging step serializes both the computation and communication and is a bottleneck when scaled to more machines.
@@
The new GreedyML algorithm brings additional parallelism to this step and can lower the memory and time required to solve the problem. We randomly partition the data among all the processors, which constitute the leaves of an accumulation tree, and then merge partial solutions at multiple levels in the tree. We prove that the GreedyML algorithm has a worst-case approximation guarantee of $1/(L+1)$ of the serial Greedy algorithm, where $L$ is the total number of accumulation levels in the accumulation tree. Using the BSP model, we analyze the time complexity of both computation and communication steps in the GreedyML  and RandGreedi algorithms, and show that the former has lower computation and communication costs than the latter.
We evaluate the parallel algorithms on the maximum $k$-set cover problem, the maximum $k$-vertex dominating set in graphs, and exemplar-based clustering (modeled by the $k$-medoid problem); all of these problems arise in data reduction or summarization. We experiment on massive data sets with millions of elements that exceed the memory constraints (a few GBs) on a single processor.
@@
We demonstrate how solutions may be computed using the parallel algorithm by organizing the accumulation tree to have more levels to adapt to the memory available on a processor. This strategy also enables us to solve for larger values of the parameter $k$ in the problems discussed above, which corresponds to the size of the solution sought. We show that the number of function evaluations on the critical path of the accumulation tree, and hence the run time, could be reduced when the parallel algorithm is employed. Also, we do not observe the deterioration in objective function values expected from the worst-case approximation ratio of the GreedyML algorithm, and the observed approximation quality of the computed solutions closely matches those obtained by the RandGreedi  algorithm on these problems.
@@ 2.1 Submodular functions
A submodular function $f$ is monotone if for every $X\subseteq Y\subseteq W$, we have $f(X)\leqslant f(Y)$. The constrained submodular maximization problem maximizes a submodular function subject to certain constraints:
@@
2.2 Lovász extension
For the analysis of our algorithm, we use the Lovász extension [20], a relaxation of submodular functions. A submodular function $f$ can be viewed as a function defined over the vertices of the unit hypercube, $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{+}$, by identifying sets $V\subseteq W$ with binary vectors of length $n=|W|$ in which the $i^{\text{th}}$ component is $1$ if $i\in V$, and 00 otherwise. The Lovász extension [20] $\widehat{f}:[0,1]^{n}\rightarrow\mathbb{R}^{+}$ is a convex extension that extends $f$ over the entire hypercube, which is given by
$\widehat{f}(x)=\mathop{\mathbb{E}}\limits_{\theta\in\mathcal{U}[0,1]}\left[\;f\left(\{i:x_{i}\geqslant\theta\}\right)\;\right].$
Here, $\theta$ is uniformly random in $[0,1]$. For any submodular function $f$, the Lovász extension $\widehat{f}$ satisfies the following properties [20]:
@@ 2.2 Lovász extension
    $\widehat{f}(1_{S})=f(S)$, for all $S\subseteq V$ where $1_{S}\in[0,1]^{n}$ is a vector containing 1 for the elements in $S$ and 00 otherwise,
@@
An $\alpha$-approximation algorithm (where $\alpha\in[0,1)$) for maximizing a submodular function $f:2^{W}\rightarrow\mathbb{R}^{+}$ subject to a hereditary constraint $\mathcal{C}$ produces a solution $S\subseteq W$ with $S\in\mathcal{C}$, satisfying $f(S)\geqslant\alpha\cdot f(S^{*})$, where $S^{*}$ is an optimal solution.
2.3 The Greedi  and RandGreedi Algorithms

### v3

# Shivaram Gopal    SM Ferdous    Hemanta Maji    Alex Pothen

@@ Abstract
We describe a parallel approximation algorithm for maximizing monotone submodular functions subject to hereditary constraints on distributed memory multiprocessors. Our work is motivated by the need to solve submodular optimization problems on massive data sets, for practical contexts such as data summarization, machine learning, and graph sparsification.
Our work builds on the randomized distributed RandGreedi algorithm, proposed by Barbosa, Ene, Nguyen, and Ward (2015). This algorithm computes a distributed solution by randomly partitioning the data among all the processors and then employing a single accumulation step in which all processors send their partial solutions to one processor. However, for large problems, the accumulation step exceeds the memory available on a processor, and the processor which performs the accumulation becomes a computational bottleneck.
Hence we propose a generalization of the RandGreedi algorithm that employs multiple accumulation steps to reduce the memory required. We analyze the approximation ratio and the time complexity of the algorithm (in the BSP model). We evaluate the new GreedyML algorithm on three classes of problems, and report results from large-scale data sets with millions of elements. The results show that the GreedyML algorithm can solve problems where the sequential Greedy and distributed RandGreedi algorithms fail due to memory constraints. For certain computationally intensive problems, the GreedyML algorithm is faster than the RandGreedi algorithm. The observed approximation quality of the solutions computed by the GreedyML algorithm closely matches those obtained by the RandGreedi  algorithm on these problems.
keywords:
Combinatorial optimization, submodular functions, distributed algorithms, approximation algorithms, data summarization.
category:
[…]
@@ 1 Introduction
We describe GreedyML, a parallel approximation algorithm for maximizing monotone submodular functions subject to hereditary constraints on distributed memory multiprocessors. GreedyML is built on an earlier distributed approximation algorithm, which has limited parallelism and higher memory requirements. Maximizing a submodular function under constraints is NP-hard, but a natural iterative Greedy algorithm exists that selects elements based on the marginal gain (defined later) and is $(1-1/\mathrm{e})\approx 0.63$-approximate for cardinality constraints and $1/2$-approximate for matroid constraints; here $\mathrm{e}$ is Euler’s number.
@@
Maximizing a submodular function (rather than a linear objective function) promotes diversity in the computed solution since at each step the algorithm encourages an element with the least properties in common with the current solution set. A broad collection of practical problems are modeled using submodular functions, including data and document summarization [21], load balancing parallel computations in quantum chemistry [7], sensor placement [5], resource allocation [26], active learning [9], interpretability of neural networks [6], influence maximization in social networks [11], diverse recommendation [4]. Surveys discussing submodular optimization formulations, algorithms, and computational experiments include Tohidi et al. [27] and Krause and Golovin [12].
@@
Our algorithm builds on the RandGreedi  framework [2], a state-of-the-art randomized distributed algorithm for monotone submodular function maximization under hereditary constraints, which has an approximation ratio half that of the Greedy algorithm. The RandGreedi algorithm randomly partitions the data among all the processors, runs the standard Greedy algorithm on each partition independently in parallel, and then executes a single accumulation step in which all processors send their partial solutions to one processor. However, this accumulation step could exceed the memory available on a processor when the memory is small relative to the size of the data, or when solutions are large. Additionally, the accumulation serializes both the computation and communication and is a bottleneck when scaled to many machines.
@@
Our GreedyML algorithm brings additional parallelism to this step and can lower the memory and running time by introducing hierarchical accumulation organized through an accumulation tree. Similar to RandGreedi, we randomly partition the data among all the processors, which constitute the leaves of the accumulation tree. We merge partial solutions at multiple levels in the tree, and the final solution is computed at the root. We prove that the GreedyML algorithm has a worst-case expected approximation guarantee of $(\alpha\>b)/(m+b)$, where $\alpha$ is the approximation guarantee for the Greedy algorithm, $b$ is the branching factor, and $m$ is the number of leaves in the accumulation tree. Using the BSP model, we also analyze the time and communication complexity of the GreedyML  and RandGreedi algorithms and show that the former has lower computation and communication costs than the latter.
@@
We evaluate the parallel algorithms on three representative and practical submodular function maximization problems: maximum $k$-set cover, maximum $k$-vertex dominating set in graphs, and exemplar-based clustering (modeled by the $k$-medoid problem). We experiment on large data sets with millions of elements that exceed the memory constraints (a few GBs) on a single processor, and demonstrate how to choose the accumulation tree to have more levels to adapt to the small memory available on a processor. This strategy also enables us to solve for larger values of the parameter $k$ in the problems discussed above, which corresponds to the size of the solution sought. We also show that the number of function evaluations on the critical path of the accumulation tree, and hence the run time, could be reduced when the parallel algorithm is employed. In most cases, we find the quality of the computed solutions by our GreedyML closely matches those obtained by the RandGreedi  algorithm on these problems despite having a worse expected approximation guarantee.
@@ 2 Background and Related Work
A submodular function $f$ is monotone if for every $X\subseteq Y\subseteq W$, we have $f(X)\leqslant f(Y)$. The constrained submodular maximization problem is defined as follows.
@@
Lovász extension
For the analysis of our algorithm, we use the Lovász extension [19], a relaxation of submodular functions. A submodular function $f$ can be viewed as a function defined over the vertices of the unit hypercube, $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{+}$, by identifying sets $V\subseteq W$ with binary vectors of length $w=|W|$ in which the $i^{\text{th}}$ component is $1$ if $i\in V$, and 00 otherwise. The Lovász extension [19] $\widehat{f}:[0,1]^{w}\rightarrow\mathbb{R}^{+}$ is a convex extension that extends $f$ over the entire hypercube and given by, $\widehat{f}(x)=\mathop{\mathbb{E}}\limits_{\theta\in\mathcal{U}[0,1]}\left[\;f\left(\{i:x_{i}\geqslant\theta\}\right)\;\right].$ Here, $\theta$ is uniformly random in $[0,1]$. The Lovász extension $\widehat{f}$ satisfies the following properties [19]:
@@
    $\widehat{f}(1_{S})=f(S)$, for all $S\subseteq V$ where $1_{S}\in[0,1]^{w}$ is a vector containing 1 for the elements in $S$ and 00 otherwise,
@@
An $\alpha$-approximation algorithm ($\alpha\in[0,1)$) for constrained submodular maximization produces a feasible solution $S\subseteq W$, satisfying $f(S)\geqslant\alpha\cdot f(S^{*})$, where $S^{*}$ is an optimal solution.
2.1 Related Work
Greedi and RandGreedi.